# Assembly Theory is an approximation to algorithmic complexity based on LZ compression that does not explain selection or evolution

### v5

@@ A.2 Assembly theory is an approximation to algorithmic complexity
One can demonstrate in Lemma 1 and Corollary 2 that the assembly index $c_{\Gamma}\left(y\right)$ (or the assembly number $A\#\left(X\right)$ ) is an approximation to algorithmic complexity $\mathbf{K}\left(y\right)$ (or, respectively, $\mathbf{K}\left(X\right)$) that in general overestimates its value as much as one wishes.
As in [4], the main avenue to achieving the following theoretical results is to construct a randomly generated program that receives a formal theory (which contains all the procedures and statistical criteria in Assembly Theory) as input. Then, it searches for an object $y$ in an assembly space with a sufficiently high assembly index $c_{\Gamma}\left(y\right)$ so as to overcome any arbitrarily chosen error margin $k$ between $c_{\Gamma}\left(y\right)$ and the algorithmic complexity $\mathbf{K}\left(y\right)$.
@@ A.2.1 Assembly number and algorithmic complexity
The reader is invited to note that the latter Lemma 3 and Corollary 4 can also be generalised for any recursive function $f$ defined on top of the assembly index and on an arbitrary observable (for which Assembly Theory, in particular, would include a formal criterion to decide whether or not function $f$ can be calculated from this observable). For example, it applies to the arbitrary function $A^{\prime\prime}\#\left(X\right)$ from Section A.3.4.
@@ A.3 Assembly theory is an encoding/compression scheme
A.3.1 Assembly index defines an encoding scheme
The following results show that the minimum rooted assembly (sub)space, from which assembly index $c_{\Gamma}\left(y\right)$ is calculated, constitutes a parameter that can be employed to encode the minimal assembling process an object $y$. More formally, Theorem 5 demonstrates that any arbitrarily chosen encoding of the set of vertices (in any arbitrary ordering) of a minimum rooted assembly space is sufficient to encode the highest-assembly-index object assembled across this space. In other words, the sheer collection of objects in a minimum rooted assembly subspace carries enough information to allow one to always pick the object with the highest assembly index value, even though the ordering of the objects in this collection can be completely arbitrary in the first place. This demonstrates that AT’s assembly index calculation method is one of the possible methods to encode objects.
@@ A.3.1 Assembly index defines an encoding scheme
Remember that ${\Gamma^{*}}_{y}$ denotes the minimum rooted assembly subspace of $\Gamma$ from which the assembly index calculates the augmented cardinality that defines the value $c_{\Gamma}\left(y\right)$. Notice that since ${\Gamma^{*}}_{y}$ is minimal, then there is a simple algorithm that can always calculate the integer value of $c_{\Gamma}\left(y\right)$ given $\mathbf{K}\left(\left<V\left({\Gamma^{*}}_{y}\right)\right>\right)$ as input. Remember that, from our definitions and conditions in Section A.1, $V\left({\Gamma^{*}}_{y}\right)\right>$ denotes an encoding of the set of vertices of ${\Gamma^{*}}_{y}$, an encoding method which is arbitrary, previously chosen, and does not depend on the choice of the assembly (sub)space or object. Now, one can exploit the computability of Assembly Theory and construct a program $p_{3}$ that receives $\mathbf{F}$ and $V\left({\Gamma^{*}}_{y}\right)\right>$ as inputs. By using $\mathbf{F}$ to decide whether or not an arbitrary $\Gamma^{\prime}$ satisfies the criteria of being an assembly subspace, program $p_{3}$ enumerates the finite set $\mathbf{\Gamma}\left(V\left({\Gamma^{*}}_{y}\right)\right)$ of all possible assembly subspaces whose set of vertices is exactly $V\left({\Gamma^{*}}_{y}\right)$. Next, using $\mathbf{F}$ to decide whether or not an arbitrary $\Gamma^{\prime}$ satisfies the criteria of being a minimum rooted assembly subspace whose longest path ends in the same last vertex of $\Gamma^{\prime}$, program $p_{3}$ searches in $\mathbf{\Gamma}\left(V\left({\Gamma^{*}}_{y}\right)\right)$ for the first assembly subspace $\Gamma^{\prime\prime}$ that is minimal. Finally, it returns the object $y^{\prime}\in V\left(\Gamma^{\prime\prime}\right)$ at which the longest rooted paths in $\Gamma^{\prime\prime}$ end (i.e., $y^{\prime}$ is the terminal vertex of these longest rooted paths). In order to achieve the desired proof, one has yet to show that the object built at the end of the longest paths in the minimum rooted assembly subspaces is unique and $y^{\prime}=y$. To this end, one can employ the minimality of ${\Gamma^{*}}_{y}$. First, suppose there is a minimum $\Gamma_{2}\in\mathbf{\Gamma}\left(V\left({\Gamma^{*}}_{y}\right)\right)$ that ends at the object $z$ such that $\Gamma_{2}\neq\Gamma^{\prime\prime}$ and $z\neq y^{\prime}$. Then, by construction of the program $p_{3}$, one would have it that $y^{\prime}$ belongs to at least one of the rooted paths in $\Gamma_{2}$. Hence, there would be a rooted path that leads to $y^{\prime}$ that is shorter than the one that leads to $z$ because by hypothesis one has it that the path that leads to $z$ is the longest one; and if there were another terminal vertex $w$ in $\Gamma_{2}$ whose rooted path has the same length of that of $z$, then $\Gamma_{2}$ would not be a minimum subspace. As a consequence, there would be a rooted assembly subspace ${\Gamma_{2}^{*}}_{y^{\prime}}\subseteq\Gamma_{2}$ whose augmented cardinality is smaller than $c_{\Gamma^{\prime\prime}}\left(y^{\prime}\right)$. However, by construction, we have it that $c_{\Gamma^{\prime\prime}}\left(y^{\prime}\right)$ is minimal, which leads to a contradiction. Therefore, since by hypothesis we have it that ${\Gamma^{*}}_{y}$ is minimal and ${\Gamma^{*}}_{y}\in\mathbf{\Gamma}\left(V\left({\Gamma^{*}}_{y}\right)\right)$, we conclude that either $\Gamma_{2}=\Gamma^{\prime\prime}={\Gamma^{*}}_{y}$ or $z=y^{\prime}=y$, and hence that, in any case, program $p_{3}$ produces $y$ as output. Finally, the proof of Equation (8) follows from basic inequalities in AIT. ∎
@@
Notice that $V\left({\Gamma^{*}}_{y}\right)\right>$ is only built to encode the membership relation for the set, and does not need to carry any a priori information about its internal structure that results in the assembly space from which it only takes the set of vertices. Only by applying a third-party decoding algorithm, such as program $p_{3}$ in the proof of Theorem 5, is that the final object $y$ becomes uniquely retrievable.
@@
One can demonstrate that from encoding one333 There might be more than one of those paths. Corollary 6 holds in any case, i.e., it holds for any arbitrarily chosen $\gamma_{y}^{\max}$ in ${\Gamma^{*}}_{y}$. of the longest444 Notice that a directly analogous version of Corollary 6 holds for $\gamma_{y}^{\min}$ instead of $\gamma_{y}^{\max}$. From encoding any of the shortest paths in a minimum rooted assembly subspace, one can always directly retrieve a minimum rooted assembly subspace, and hence the terminal vertex that corresponds to the object to be assembled. paths in a minimum rooted assembly subspace, one can always retrieve a minimum rooted assembly subspace, and hence the terminal vertex that corresponds to the object to be assembled:
@@
holds, where $\gamma\right>$ is any arbitrarily fixed encoding of a sequence $\gamma$ composed by contiguous oriented edges together with the starting vertex/object in the basis set $B_{\Gamma}\subseteq B_{\mathcal{S}}$, sequence which defines the path $\gamma$.
@@
Let $p_{3}$ be the program used in the proof of Theorem 5. Let $p_{4}$ be a program that receives $\mathbf{F}$, program $p_{3}$, and $\gamma_{y}^{\max}\right>$ as inputs. Then, by using $\mathbf{F}$ to decide whether or not an arbitrary $\Gamma^{\prime}$ satisfies the criteria of being a minimum rooted assembly subspace, enumerates all possible minimum rooted assembly subspaces that have $\gamma_{y}^{\max}$ as one of its longest rooted paths. In the next step, program $p_{4}$ picks the first of these possible minimum rooted assembly subspaces and encodes it in the form $V\left({\Gamma^{*}}_{y}\right)\right>$. Finally, program $p_{4}$ returns the output of program $p_{3}$ with $\mathbf{F}$ and $V\left({\Gamma^{*}}_{y}\right)\right>$ as inputs. The proof of Equation (9) follows from Equation (8) and basic inequalities in AIT. ∎
A sequence of edges in each of these longest (or shortest) paths constitutes a way to parse the object (e.g., a string), i.e., to decompose it into contiguous subsequences. Each distinct assembling path in ${\Gamma^{*}}_{y}$ correspond to a distinct way to parse the object $y$. This idea will become clearer in the next Sections A.3.2 and A.3.3. It also constitutes the compression scheme underlying LZ algorithms.
A.3.2 Assembly index defines a grammar of compression that is bounded by LZ compression and can at most approximate the Shannon Entropy rate
@@
As presented in Section A.1, recall that the assembly index $c_{\Gamma}\left(x\right)$ of an object $x$ is defined as the size of the smallest acyclic quiver (directed multigraph) with edge-labellings from the very set of vertices555 And also by adding the property described in Item $6$ in [53, Definition 11] of assembly spaces. , a quiver which is called the assembly (sub)space, denoted by $\left(\Gamma,\phi\right)$, or just $\Gamma$ where the edge-labellings given by $\phi$ are not relevant. The “tree”-like structure of the assembly space maps the generation of an object from a set of basis objects in $B_{\mathcal{S}}$, which is a finite set and remains constant for any object assembled from combinations of the basis objects. Each vertex represents an object, and the directed edge relation specifies the generation hierarchy (i.e., the label of the oriented edge) of the object to be combined with the previous object (i.e., the source vertex to which the respective outgoing labelled edges are connected) in order to assemble a new object (i.e., the target vertex to which the respective incoming labelled edges are connected).
@@
It is straightforward to show the existence of such a CFG that encodes all the objects generated by an assembly space. In the following, we will investigate it further to demonstrate that for any ${\Gamma^{*}}_{x}$ (i.e., a minimal rooted assembly subspace) that generates a string $x$, there exists a grammar $G$ that encodes this ${\Gamma^{*}}_{x}$, including not only the final object $x$ to be assembled, but encoding the whole assembling process rooted on the basis objects. In other words, there is at least one CFG that generates $x$ and manifests a correspondence between: the basis objects of both the grammar (terminal symbols) and the assembly subspace; non-basis objects and vertices of ${\Gamma^{*}}_{y}$ with the sets of non-terminal symbols; and edges of ${\Gamma^{*}}_{y}$ with derivation rules of the grammar.
@@
Let ${\Gamma^{*}}_{x}$ generates $x$. First, we define the set of terminal symbols in $G$ as the set of basis objects in ${\Gamma^{*}}_{x}$ plus the final object $x$ to be assembled, that is, $\Sigma_{G}=B_{\Gamma}\cup\left\{x\right\}$. Then, let $V_{G}\setminus\Sigma_{G}=V\left({\Gamma^{*}}_{x}\right)\setminus\left\{x\right\}$ be the set of non-terminal symbols. Now, we construct the set $R$ of derivation rules of the CFG in the following way: we start by adding $S\rightarrow x$ for each basis object $x$ that belongs to $V_{G}\setminus\Sigma_{G}$ and also for each basis object that belongs to $\Sigma_{G}$ to the set of derivation rules666 Thus, there will be $2\left|B_{\Gamma}\right|$ rules in the form $S\rightarrow x$.; then, for each $e_{i}\in E\left({\Gamma^{*}}_{x}\right)$ such that $s_{\Gamma}(e_{i})=x_{i}$ and $t_{\Gamma}(e_{i})=y_{i}$, we construct a corresponding new rule, either $x_{i}\rightarrow x_{i}z_{i}$ or $x_{i}\rightarrow z_{i}x_{i}$, where $z_{i}\in V\left({\Gamma^{*}}_{x}\right)$ and $\phi\left(e_{i}\right)=z_{i}$, whichever concatenation of combinations builds the object/vertex $y_{i}$ (i.e., either $y_{i}=x_{i}z_{i}$ or $y_{i}=z_{i}x_{i}$). This is possible since each “assembly step” (represented by the respective edge) can only combine two objects: one object is the source vertex of the edge; and the other object is the vertex that is the label of this edge. If $y_{i}=x_{i}z_{i}\neq x$ (or $y_{i}=z_{i}x_{i}\neq x$), then we add $y_{i}$ to the set of non-terminal symbols $V_{G}\setminus\Sigma_{G}$. Otherwise, we add $y_{i}=x$ to the set of terminal symbols of $G$. From this construction of the set $R$, one will obtain that
@@
The grammar defined in Theorem 7 demonstrates that any assembled object can be encoded into a Context-Free Grammar (CFG) in Chomsky Normal Form (CNF) that captures the same information regarding the relationship of how an object is constructed from other objects. The assembly index is defined as the number of non-basis vertices, while the size of the grammar is defined as the number of derivation rules, which is equivalent to the number of nonterminals if there are no useless rules [15]. Let’s denote this size by $|G|$. In order to encode the object $x$ assembled according to ${\Gamma^{*}}_{x}$, there is no necessity for having two or more edges linking the same pair of vertices in an assembly space. Thus, for completeness’ sake, we will denote by $G^{*}$ the grammar with any useless rules/edges removed. We will show in Corollary 8 that there is a simple linear relationship between the assembly index $c_{\Gamma}(x)$ and the size of the grammar $|G^{*}|$, except for the size of the basis set $B_{\Gamma}$ which remains constant for any object assembled upon combinations of basis objects.
@@
Let $x$ be a string/object assembled according to ${\Gamma^{*}}_{x}$ and $G$ be the CFG in Theorem 7. Then,
@@
Corollary 8 proves that the assembly index for strings is equivalent to the compression size of a compressing grammar, demonstrating that the notion of the minimum “amount of assembling steps” (that the minimum rooted assembly subspace defines) is exactly the same as a (minimal) CFG. For each of the grammars constructed in Theorem 7, the final object $x$ is unique, and for each distinct assembling path leading to $x$, there is a (unique) correspondent sequence of rules in $R$ that also results in the object $x$.
@@
A.3.3 Assembly index is a version of LZ compression
In Section A.3, we have demonstrated that either the set of vertices of the minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ or at least one of the longest (or shortest) rooted paths in this ${\Gamma^{*}}_{y}$ is sufficient to encode the object $y$. These results employ general algorithmic methods so that they are independent of the choice of the programming language and encoding-decoding schemes. In Theorem 7, we have demonstrated that every minimum rooted assembly (sub)space corresponds to at least one context-free grammar (CFG), which reveals a way one can compress an assembly space (e.g., the minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ that generates the object $y$) into the minimum set of rules of a CFG. These results together indicate the existence of some LZ-based method to encode both the object and its assembling process. Indeed, the proof of the following Theorem 9 constructs such a LZ compression scheme.
For traditional LZ schemes, the encoded/compressed form of a string comprises a sequence of encoded tuples (or factors) of finite length. Typically, each one of these tuples contains at least one pointer (e.g., in the case of LZ78787878 and LZW, called indexes) that refers to another tuple or to an entry in a dictionary. Also, it can contain information about some elements of the basic alphabet (e.g., in the case of LZ78787878 and LZW, it contains the codeword for only one symbol to be concatenated at the end of the element the pointer is referring to). It is trivial to find examples of objects whose parsing according to the LZ78787878 differs from the parsing according to the LZ77777777, or that the one according to the LZMA differs from LZW, and so on. The LZ family tree of compression schemes includes many distinct schemes, including LZ77, LZ78, LZW, LZSS, and LZMA [62]. All of these compression methods are defined by schemes that resort to a (static or dynamic) dictionary containing tokens, indexes, pointers, or basic symbols from which the decoder or decompressor can retrieve the original raw data from the received encoded/compressed form according to the respective LZ scheme [62]. Theorem 9 demonstrates that the minimum rooted assembly (sub)space from which the assembly index is calculated is a compression scheme that belongs to this LZ family of compression algorithms.
On the one hand, the pointers in traditional LZ compression schemes, such as LZ77, LZ78 or LZW, refer to encoded tuples that have occurred before as recorded by the associated dictionary. On the other hand, for Assembly Theory, the pointers refer to other tuples that might have not occurred before in an assembling path, but are part of the minimum rooted assembly subspace ${\Gamma^{*}}_{x}$, that is, part of a “tree”-like structure that needs to be encoded in the dictionary, which is only built or queried during the encoding/decoding process.

### v6

@@ A.2 Assembly theory is an approximation to algorithmic complexity
One can demonstrate in Lemma 1 and Corollary 2 that the assembly index $c_{\Gamma}\left(y\right)$ (or the assembly number $A\#\left(X\right)$ ) is an approximation to algorithmic complexity $\mathbf{K}\left(y\right)$ (or, respectively, $\mathbf{K}\left(X\right)$) that in general overestimates its value as much as desired.
As in [4], the main route to achieving the following theoretical results is to construct a randomly generated program that receives a formal theory (which contains all the procedures and statistical criteria in Assembly Theory) as input. Then, it searches for an object $y$ in an assembly space with a sufficiently high assembly index $c_{\Gamma}\left(y\right)$ so as to overcome any arbitrarily chosen error margin $k$ between $c_{\Gamma}\left(y\right)$ and the algorithmic complexity $\mathbf{K}\left(y\right)$.
@@ A.2.1 Assembly number and algorithmic complexity
The reader is invited to note that the latter Lemma 3 and Corollary 4 can also be generalised for any recursive function $f$ defined on the basis of the assembly index and on an arbitrary observable (for which Assembly Theory, in particular, would include a formal criterion to decide whether or not function $f$ can be calculated from this observable). For example, it applies to the arbitrary function $A^{\prime\prime}\#\left(X\right)$ from Section A.3.4.
@@ A.3 Assembly theory is an encoding/compression scheme
A.3.1 The assembly index defines an encoding scheme
The following results show that the minimum rooted assembly (sub)space, from which assembly index $c_{\Gamma}\left(y\right)$ is calculated, constitutes a parameter that can be employed to encode the minimal assembling process of an object $y$. More formally, Theorem 5 demonstrates that any arbitrarily chosen encoding of the set of vertices (in any arbitrary ordering) of a minimum rooted assembly space is sufficient to encode the highest-assembly-index object assembled across this space. In other words, the sheer number of objects in a minimum rooted assembly subspace carries enough information to allow one to always pick the object with the highest assembly index value, even though the ordering of the objects in this collection may be completely arbitrary in the first place. This demonstrates that AT’s assembly index calculation method is one of the possible methods for encoding objects.
@@ A.3.1 The assembly index defines an encoding scheme
Remember that ${\Gamma^{*}}_{y}$ denotes the minimum rooted assembly subspace of $\Gamma$ from which the assembly index calculates the augmented cardinality that defines the value $c_{\Gamma}\left(y\right)$. Notice that since ${\Gamma^{*}}_{y}$ is minimal, then there is a simple algorithm that can always calculate the integer value of $c_{\Gamma}\left(y\right)$ given $\mathbf{K}\left(\left<V\left({\Gamma^{*}}_{y}\right)\right>\right)$ as input. Remember that, from our definitions and conditions in Section A.1, $V\left({\Gamma^{*}}_{y}\right)\right>$ denotes an encoding of the set of vertices of ${\Gamma^{*}}_{y}$, an encoding method which is arbitrary, previously chosen, and does not depend on the choice of the assembly (sub)space or object. Now, one can exploit the computability of Assembly Theory and construct a program $p_{3}$ that receives $\mathbf{F}$ and $V\left({\Gamma^{*}}_{y}\right)\right>$ as inputs. By using $\mathbf{F}$ to decide whether or not an arbitrary $\Gamma^{\prime}$ satisfies the criteria of being an assembly subspace, program $p_{3}$ enumerates the finite set $\mathbf{\Gamma}\left(V\left({\Gamma^{*}}_{y}\right)\right)$ of all possible assembly subspaces whose set of vertices is exactly $V\left({\Gamma^{*}}_{y}\right)$. Next, using $\mathbf{F}$ to decide whether or not an arbitrary $\Gamma^{\prime}$ satisfies the criteria of being a minimum rooted assembly subspace whose longest path ends in the same last vertex of $\Gamma^{\prime}$, program $p_{3}$ searches in $\mathbf{\Gamma}\left(V\left({\Gamma^{*}}_{y}\right)\right)$ for the first assembly subspace $\Gamma^{\prime\prime}$ that is minimal. Finally, it returns the object $y^{\prime}\in V\left(\Gamma^{\prime\prime}\right)$ at which the longest rooted paths in $\Gamma^{\prime\prime}$ end (i.e., $y^{\prime}$ is the terminal vertex of these longest rooted paths). In order to achieve the desired proof, one has yet to show that the object built at the end of the longest paths in the minimum rooted assembly subspaces is unique and that $y^{\prime}=y$. To this end, one can employ the minimality of ${\Gamma^{*}}_{y}$. First, suppose there is a minimum $\Gamma_{2}\in\mathbf{\Gamma}\left(V\left({\Gamma^{*}}_{y}\right)\right)$ that ends at the object $z$ such that $\Gamma_{2}\neq\Gamma^{\prime\prime}$ and $z\neq y^{\prime}$. Then, by construction of the program $p_{3}$, one would have it that $y^{\prime}$ belongs to at least one of the rooted paths in $\Gamma_{2}$. Hence, there would be a rooted path that leads to $y^{\prime}$ that is shorter than the one that leads to $z$ because by hypothesis one has it that the path that leads to $z$ is the longest one; and if there were another terminal vertex $w$ in $\Gamma_{2}$ whose rooted path has the same length as that of $z$, then $\Gamma_{2}$ would not be a minimum subspace. As a consequence, there would be a rooted assembly subspace ${\Gamma_{2}^{*}}_{y^{\prime}}\subseteq\Gamma_{2}$ whose augmented cardinality is smaller than $c_{\Gamma^{\prime\prime}}\left(y^{\prime}\right)$. However, by construction, we have it that $c_{\Gamma^{\prime\prime}}\left(y^{\prime}\right)$ is minimal, which leads to a contradiction. Therefore, since by hypothesis we have it that ${\Gamma^{*}}_{y}$ is minimal and ${\Gamma^{*}}_{y}\in\mathbf{\Gamma}\left(V\left({\Gamma^{*}}_{y}\right)\right)$, we conclude that either $\Gamma_{2}=\Gamma^{\prime\prime}={\Gamma^{*}}_{y}$ or $z=y^{\prime}=y$, and hence that, in any case, program $p_{3}$ produces $y$ as output. Finally, the proof of Equation (8) follows from basic inequalities in AIT. ∎
@@
Notice that $V\left({\Gamma^{*}}_{y}\right)\right>$ is only built to encode the membership relation for the set, and does not need to carry any a priori information about its internal structure that results in the assembly space from which it only takes the set of vertices. Only by applying a third-party decoding algorithm, such as program $p_{3}$ in the proof of Theorem 5, does the final object $y$ become uniquely retrievable.
@@
One can demonstrate that from encoding one333 There might be more than one of these paths. Corollary 6 holds in any case, i.e., it holds for any arbitrarily chosen $\gamma_{y}^{\max}$ in ${\Gamma^{*}}_{y}$. of the longest444 Notice that a directly analogous version of Corollary 6 holds for $\gamma_{y}^{\min}$ instead of $\gamma_{y}^{\max}$. From encoding any of the shortest paths in a minimum rooted assembly subspace, one can always directly retrieve a minimum rooted assembly subspace, and hence the terminal vertex that corresponds to the object to be assembled. paths in a minimum rooted assembly subspace, one can always retrieve a minimum rooted assembly subspace, and hence the terminal vertex that corresponds to the object to be assembled:
@@
holds, where $\gamma\right>$ is any arbitrarily fixed encoding of a sequence $\gamma$ composed of contiguous oriented edges together with the starting vertex/object in the basis set $B_{\Gamma}\subseteq B_{\mathcal{S}}$, the sequence which defines the path $\gamma$.
@@
Let $p_{3}$ be the program used in the proof of Theorem 5. Let $p_{4}$ be a program that receives $\mathbf{F}$, program $p_{3}$, and $\gamma_{y}^{\max}\right>$ as inputs. Then, by using $\mathbf{F}$ to decide whether or not an arbitrary $\Gamma^{\prime}$ satisfies the criteria of being a minimum rooted assembly subspace, it enumerates all possible minimum rooted assembly subspaces that have $\gamma_{y}^{\max}$ as one of its longest rooted paths. In the next step, program $p_{4}$ picks the first of these possible minimum rooted assembly subspaces and encodes it in the form $V\left({\Gamma^{*}}_{y}\right)\right>$. Finally, program $p_{4}$ returns the output of program $p_{3}$ with $\mathbf{F}$ and $V\left({\Gamma^{*}}_{y}\right)\right>$ as inputs. The proof of Equation (9) follows from Equation (8) and basic inequalities in AIT. ∎
A sequence of edges in each of these longest (or shortest) paths constitutes a way to parse the object (e.g., a string), i.e., to decompose it into contiguous subsequences. Each distinct assembling path in ${\Gamma^{*}}_{y}$ corresponds to a distinct way to parse the object $y$. This idea will become clearer in the next Sections A.3.2 and A.3.3. It also constitutes the compression scheme underlying LZ algorithms.
A.3.2 The assembly index defines a grammar of compression that is bounded by LZ compression and can at most approximate the Shannon Entropy rate
@@
As presented in Section A.1, recall that the assembly index $c_{\Gamma}\left(x\right)$ of an object $x$ is defined as the size of the smallest acyclic quiver (directed multigraph) with edge-labellings from its own set of vertices555 And also by adding the property described in Item $6$ in [54, Definition 11] of assembly spaces. , a quiver which is called the assembly (sub)space, denoted by $\left(\Gamma,\phi\right)$, or just $\Gamma$ where the edge-labellings given by $\phi$ are not relevant. The “tree”-like structure of the assembly space maps the generation of an object from a set of basis objects in $B_{\mathcal{S}}$, which is a finite set and remains constant for any object assembled from combinations of the basis objects. Each vertex represents an object, and the directed edge relation specifies the generation hierarchy (i.e., the label of the oriented edge) of the object to be combined with the previous object (i.e., the source vertex to which the respective outgoing labelled edges are connected) in order to assemble a new object (i.e., the target vertex to which the respective incoming labelled edges are connected).
@@
It is straightforward to show the existence of such a CFG that encodes all the objects generated by an assembly space. In the following, we will investigate it further to demonstrate that for any ${\Gamma^{*}}_{x}$ (i.e., a minimal rooted assembly subspace) that generates a string $x$, there exists a grammar $G$ that encodes this ${\Gamma^{*}}_{x}$, encoding not only the final object $x$ to be assembled, but the whole assembling process rooted in the basis objects. In other words, there is at least one CFG that generates $x$ and manifests a correspondence between: the basis objects of both the grammar (terminal symbols) and the assembly subspace; non-basis objects and vertices of ${\Gamma^{*}}_{y}$ with the sets of non-terminal symbols; and edges of ${\Gamma^{*}}_{y}$ with derivation rules of the grammar.
@@
Let ${\Gamma^{*}}_{x}$ generate $x$. First, we define the set of terminal symbols in $G$ as the set of basis objects in ${\Gamma^{*}}_{x}$ plus the final object $x$ to be assembled, that is, $\Sigma_{G}=B_{\Gamma}\cup\left\{x\right\}$. Then, let $V_{G}\setminus\Sigma_{G}=V\left({\Gamma^{*}}_{x}\right)\setminus\left\{x\right\}$ be the set of non-terminal symbols. Now, we construct the set $R$ of derivation rules of the CFG in the following way: we start by adding $S\rightarrow x$ for each basis object $x$ that belongs to $V_{G}\setminus\Sigma_{G}$ and also for each basis object that belongs to $\Sigma_{G}$ to the set of derivation rules666 Thus, there will be $2\left|B_{\Gamma}\right|$ rules in the form $S\rightarrow x$.; then, for each $e_{i}\in E\left({\Gamma^{*}}_{x}\right)$ such that $s_{\Gamma}(e_{i})=x_{i}$ and $t_{\Gamma}(e_{i})=y_{i}$, we construct a corresponding new rule, either $x_{i}\rightarrow x_{i}z_{i}$ or $x_{i}\rightarrow z_{i}x_{i}$, where $z_{i}\in V\left({\Gamma^{*}}_{x}\right)$ and $\phi\left(e_{i}\right)=z_{i}$, whichever concatenation of combinations builds the object/vertex $y_{i}$ (i.e., either $y_{i}=x_{i}z_{i}$ or $y_{i}=z_{i}x_{i}$). This is possible since each “assembly step” (represented by the respective edge) can only combine two objects: one object is the source vertex of the edge; and the other object is the vertex that is the label of this edge. If $y_{i}=x_{i}z_{i}\neq x$ (or $y_{i}=z_{i}x_{i}\neq x$), then we add $y_{i}$ to the set of non-terminal symbols $V_{G}\setminus\Sigma_{G}$. Otherwise, we add $y_{i}=x$ to the set of terminal symbols of $G$. From this construction of the set $R$, one will obtain that
@@
The grammar defined in Theorem 7 demonstrates that any assembled object can be encoded into a Context-Free Grammar (CFG) in Chomsky Normal Form (CNF) that captures the same information regarding how an object is constructed from other objects. The assembly index is defined as the number of non-basis vertices, while the size of the grammar is defined as the number of derivation rules, which is equivalent to the number of nonterminals if there are no useless rules [15]. Let’s denote this size by $|G|$. In order to encode the object $x$ assembled according to ${\Gamma^{*}}_{x}$, there is no necessity for having two or more edges linking the same pair of vertices in an assembly space. Thus, for completeness’ sake, we will denote by $G^{*}$ the grammar with any useless rules/edges removed. We will show in Corollary 8 that there is a simple linear relationship between the assembly index $c_{\Gamma}(x)$ and the size of the grammar $|G^{*}|$, except for the size of the basis set $B_{\Gamma}$ which remains constant for any object assembled upon combinations of basis objects.
@@
Let $x$ be a string/object assembled according to ${\Gamma^{*}}_{x}$ and let $G$ be the CFG in Theorem 7. Then,
@@
Corollary 8 proves that the assembly index for strings is equivalent to the compression size of a compressing grammar, demonstrating that the notion of the minimum “amount of assembling steps” (that the minimum rooted assembly subspace defines) is exactly the same as a (minimal) CFG. For each of the grammars constructed in Theorem 7, the final object $x$ is unique, and for each distinct assembling path leading to $x$, there is a (unique) corresponding sequence of rules in $R$ that also results in the object $x$.
@@
A.3.3 The assembly index is a version of LZ compression
In Section A.3, we have demonstrated that either the set of vertices of the minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ or at least one of the longest (or shortest) rooted paths in this ${\Gamma^{*}}_{y}$ is sufficient to encode the object $y$. These results employ general algorithmic methods so that they are independent of the choice of programming language and encoding-decoding schemes. In Theorem 7, we have demonstrated that every minimum rooted assembly (sub)space corresponds to at least one context-free grammar (CFG), which reveals a way one can compress an assembly space (e.g., the minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ that generates the object $y$) into the minimum set of rules of a CFG. These results together indicate the existence of some LZ-based method to encode both the object and its assembling process. Indeed, the proof of the following Theorem 9 constructs such a LZ compression scheme.
For traditional LZ schemes, the encoded/compressed form of a string comprises a sequence of encoded tuples (or factors) of finite length. Typically, each one of these tuples contains at least one pointer (e.g., in the case of LZ78787878 and LZW, called indexes) that refers to another tuple or to an entry in a dictionary. Also, it can contain information about some elements of the basic alphabet (e.g., in the case of LZ78787878 and LZW it contains the codeword for only one symbol to be concatenated at the end of the element the pointer is referring to). It is trivial to find examples of objects whose parsing according to the LZ78787878 differs from the parsing according to the LZ77777777, or that the one according to the LZMA differs from LZW, and so on. The LZ family tree of compression schemes includes many distinct schemes, including LZ77, LZ78, LZW, LZSS, and LZMA [63]. All of these compression methods are defined by schemes that resort to a (static or dynamic) dictionary containing tokens, indexes, pointers, or basic symbols from which the decoder or decompressor can retrieve the original raw data from the received encoded/compressed form according to the respective LZ scheme [63]. Theorem 9 demonstrates that the minimum rooted assembly (sub)space from which the assembly index is calculated is a compression scheme that belongs to this LZ family of compression algorithms.
On the one hand, the pointers in traditional LZ compression schemes, such as LZ77, LZ78 or LZW, refer to encoded tuples that have occurred before as recorded by the associated dictionary. On the other hand, for Assembly Theory, the pointers refer to other tuples that might not have occurred before in an assembling path, but are part of the minimum rooted assembly subspace ${\Gamma^{*}}_{x}$, that is, part of a “tree”-like structure that needs to be encoded in the dictionary, which is only built or queried during the encoding/decoding process.